# Bigger, Regularized, Optimistic: scaling for compute and sample-efficient continuous control

### v1

@@ Appendix E Hyperparameters
Hyperparameters of BRO and other baselines are listed in Table 6. BRO (Fast) shares the same parameters as BRO except replay ratio $2$ which significantly speeds the algorithm without sacrificing performance that much. BRO features the BRONet architecture and resets of all parameters done every $250k$ steps until $1M$ steps with additional resets at steps $15k$ and $50k$.
The selection of hyperparameters for BRO was based on the values reported in the main building blocks of BRO and extensive experimentation coupled with ablations studies. We discuss hyperparameter selection for baselines in Appendix 21.
@@
Table 6: Hyperparameter values for actor-critic agents used in the experiments. We did not rerun TD-MPC2 (Hansen et al., 2023) and report the official results.
@@
Appendix F Baselines Details
We compare BRO against official and widely used implementations of CrossQ, SAC, SR-SAC, TD3 and TD-MPC2 with open source repositories listed in Table 7. As the official results do not cover all $40$ benchmarking tasks, we ran the baselines independently (except TD-MPC2, where all official results were available).
SAC and TD3 are commonly used baselines; therefore, their hyperparameters vary across different implementations. In order to account for this fact, we ran $2$ versions of these baselines: tuned and original. If not specified otherwise, we report the results of the tuned versions with hyperparameters in Table 6. The original versions of SAC and TD3 both feature a replay ratio of 1 and in the case of SAC, target entropy ($\mathcal{H}^{*}$) equal to the action space dimension $|\mathcal{A}|$. The performance of both variants of the implementations can be observed in Figure 22.
As other baselines were developed and tested on only a subset of our 40 selected tasks, we observed that achieving similar performance on new tasks was challenging. This can be especially observed in the case of CrossQ, which is a state-of-the-art algorithm on selected tasks from OpenAI Gym (Brockman et al., 2016), but as it was tested only on a fraction of DeepMind Control Suite tasks, its performance does not transfer to our selection of tasks.
The RL research community also needs to be careful as changes in environment tasks wrappers can have significant effects on the performance. Originally, CrossQ authors tested their agent on DeepMind Control Suite using Shimmy (Tai et al., ) contrary to other agents that use the original codebase (Tassa et al., 2018). If not stated otherwise, all CrossQ results presented in this paper were run using DMC wrappers, but in order to test BRO thoroughly, we also ran CrossQ on 10 DMC tasks using Shimmy and present the results in Figure 15 (the remaining 5 tasks crashed, due to a bug in the official codebase which is also a warning to researchers benchmarking the agents on a very small fraction of a benchmark).
Figure 15: Results of $10$ tasks from DeepMind Control Suite for BRO and CrossQ that was run using Shimmy or DMC wrappers. Remaining $5$ tasks are not present as the official CrossQ code crashes due to bugs. We contacted the authors regarding this issue. Regardless of the used environment variants, BRO outperforms CrossQ.
@@ Appendix F Baselines Details
Table 7: Links to the repositories of the used baselines. All repositories are distributed under MIT license.
Appendix G Experimental Results
We present the aggregated performance of BRO compared to other baselines at Dog & Humanoid tasks, DeepMind Control Suite, Metaworld and MyoSuite in Figure 18 together with summarized performance results at $100k$, $200k$, $500k$ and $1M$ steps in Table 8. The performance on each of the $40$ individual tasks in shown in Figures 19, 20, 21.
(a) DeepMind Control (DMC)
(b) MetaWorld (MW)
(c) MyoSuite (MS)
@@ Appendix G Experimental Results
Figure 17: We compare BRO against BRO without target network. $10$ seeds per task, $1M$ environment steps.
@@
Table 8: Summary of IQM results of BRO and other agents evaluated on $40$ tasks from DeepMind Control Suite, Metaworld and MyoSuite achieved at $100k$, $200k$, $500k$ and $1M$ steps. BRO achieves better results than other state-of-the art agents (both model-based and model-free) while featuring great sample efficiency.
@@
Appendix H Comparison between variants of SAC and TD3
Comparison between $2$ variants of SAC and TD3 (Original and Tuned) is present in Figure 22. Tuned versions feature a higher value of replay ratio ($2$ instead of $1$) than original and lower target entropy in case of SAC ($|\mathcal{A}|/2$ instead of $|\mathcal{A}|$). BRO greatly outperforms both variants of these baselines.

### v2

@@ Appendix D Hyperparameters
Hyperparameters of BRO and other baselines are listed in Table 5. BRO (Fast) shares the same parameters as BRO except replay ratio $2$ which significantly speeds the algorithm without sacrificing performance that much. BRO features the BRONet architecture and resets of all parameters done every $250k$ steps until $1M$ steps with additional resets at steps $15k$ and $50k$. The selection of hyperparameters for BRO was based on the values reported in the main building blocks of BRO and extensive experimentation coupled with ablations studies.
@@
Table 5: Hyperparameter values for actor-critic agents used in the experiments.
@@
We compare BRO against official and widely used implementations of CrossQ, SAC, SR-SAC, TD3 and TD-MPC2 with open source repositories listed in Table 6. As the official results do not cover all $40$ benchmarking tasks, we ran the baselines independently (except TD-MPC2, where all official results were available). SAC and TD3 are commonly used baselines; therefore, their hyperparameters vary across different implementations. To account for this fact, we ran $2$ versions of these baselines: tuned and original. If not specified otherwise, we report the results of the tuned versions with hyperparameters in Table 5. The original versions of SAC and TD3 both feature a replay ratio of 1 and in the case of SAC, target entropy ($\mathcal{H}^{*}$) equal to the action space dimension $|\mathcal{A}|$. The performance of both variants of the implementations can be observed in Figure 24.
Table 6: Links to the repositories of the used baselines. All are distributed under MIT license.
@@
As other baselines were developed and tested on only a subset of our 40 selected tasks, we observed that achieving similar performance on new tasks was challenging. This can be especially observed in the case of CrossQ, which is a state-of-the-art algorithm on selected tasks from OpenAI Gym (Brockman et al., 2016), but as it was tested only on a fraction of DeepMind Control Suite tasks, its performance does not transfer to our selection of tasks. Originally, CrossQ authors tested their agent on DeepMind Control Suite using Shimmy (Tai et al., ) contrary to other agents that use the original codebase (Tassa et al., 2018). We run CrossQ using the DMC wrappers (D’Oro et al., 2022). Comparison between $2$ variants of SAC and TD3 (Original and Tuned) is presented in Figure 24. Tuned versions feature a higher value of replay ratio ($2$ instead of $1$) than the original and lower target entropy in the case of SAC ($|\mathcal{A}|/2$ instead of $|\mathcal{A}|$).
Table 7: Description of the considered model sizes.
Size Number of block Hidden Size
$0.55M$ 1 BroNet block hidden size of $128$
$1.05M$ 1 BroNet block hidden size of $256$
$2.83M$ 1 BroNet block hidden size of $512$
$4.92M$ 2 BroNet blocks hidden size of $512$
$26.31M$ 3 BroNet blocks hidden size of $1024$
Appendix E Additional Experiments
E.1 Scaling and Time Experiments
The execution time was measured for all agents for each of the $40$ tasks averaged over $2$ random seeds. We ran each agent for $105k$ steps with initial $5k$ exploration steps, $100k$ training steps, and $1$ evaluation. We benchmark all $25$ variants of BRO with $5$ different model sizes and $5$ values of replay ratio. Figure 12 different algorithms performance compared to execution time. Experiments were conducted on an NVIDIA A100 GPU with 10GB of RAM and 8 CPU cores of AMD EPYC 7742 processor. All tasks were run separately so the agents could use all resources independently.
Figure 12: Scatterplot of the performance of agents plotted against normalized execution time.
Increasing the model size and replay ratio both improve the performance. However, the former is more efficient in terms of execution time due to GPU parallelism. For example, the largest BRO variant ($26.31M$ parameters) with replay ratio $5$ has similar execution time as the smallest one ($0.55M$ parameters) with replay ratio $15$, but the performance is much greater (Figures 13).
Figure 13: Heatmaps of execution time and IQM performance across $40$ tasks of $25$ BRO variants with various model sizes and replay ratio values. Black lines connect the same interpolated values.
E.2 Additional BroNet Analysis
We examine various architectural blueprints on $5$ DMC and $5$ MetaWorld environments (see Table 3), each with over $10$ seeds per task. Our starting point was the transformer-based design by Bjorck et al. (2021), termed Spectral. This architecture incorporates recent transformer advancements, moving Layer Norm to the beginning of the residual block to prevent vanishing gradients in deep networks. While Spectral performs better than the vanilla MLP, its performance on the DMC benchmark, particularly the Dog environment, is weaker. This aligns with findings from Nauman et al. (2024); Hussing et al. (2024), indicating that Layer Norm is crucial for stability in such complex tasks. To analyze the importance of layer normalization in the BroNet architecture, we replaced spectral norms with Layer Norms in the residual blocks, resulting in the BRO wo first LN architecture (Figure 5). This modification improves performance but still lags behind the full BRO architecture. Furthermore, we examine a simple MLP architecture with Layer Norm before each activation function. Since BRO consists of two residual blocks, we compare it with a 5-layer model, (Dense + LN) x 5. Figure 14 shows that Layer Norm after each Dense layer is effective, and in aggregated IQM, this model is comparable to BRO. However, skip connections in BRO are beneficial for managing complex environments like Dog. In conclusion, BroNet architecture uses Layer Norm and residual blocks for superior robustness and performance in challenging tasks.
Figure 14: Comparison of five architecture designs across different environments: The top plot shows results on 5 DMC and 5 MetaWorld environments, the middle plot focuses on the 5 DMC environments, and the bottom plot highlights the Dog Trot environment. BRO and Spectral architectures each consist of 2 residual blocks. (Dense + LN) x 5 represents standard MLP networks with 5 hidden layers, each incorporating Layer Norm before activation. Lastly, BRO wo first LN refers to the BRO architecture without Layer Norm in the first Dense block, before the residual connection.
E.3 Additional analysis
Batch sizes
We ablate of the minibatch size impact on BRO and BRO (Fast) performance across different benchmarks is depicted in Figure 15. The figure shows that using half or even a quarter of the original minibatch size ($256$) does not significantly hurt BRO’s performance.
Figure 15: Performance of BRO and BRO (Fast) with different minibatch sizes for: D&H (Dogs and Humanoid), DMC (DeepMind Control), MW (MetaWorld), and MS (MyoSuite).
Target network
We investigate the performance benefits stemming from using a target network with the BRO agent. We present these results in Figure 16. Interestingly, we observe that the target network yields performance improvements only in specific environments.
@@ E.3 Additional analysis
Figure 16: We compare BRO against BRO without target network. $10$ seeds per task, $1M$ steps.
More baselines
To evaluate BRO performance beyond maximum entropy objectives, we tested BRO and BRO (Fast) with a TD3 backbone. BRO with a SAC backbone slightly outperformed TD3, though TD3 remains a viable option. Furthermore, we compare BRO performance to three additional baselines. These results are presented in Figure 17.
Figure 17: We run BRO and BRO (Fast) with a TD3 backbone (BROTD3). 10 seeds.
Longer training
We expanded BRO training beyond 1M environment steps, although in a single-task setup. We trained BRO and BRO (Fast) for 3M and 5M steps respectively on 7 Dog and Humanoid tasks and compared them to TD-MPC2 and SR-SAC. BRO significantly outperforms these baselines and notably almost solves the Dog Run tasks at 5M steps (achieving over 80% of possible returns). We show the 3M results in Figure 18.
Figure 18: We run BRO on complex tasks for 3M steps. 5 seeds.
Image-based tasks
To analyze the impact of BroNet on image-based RL tasks, we experiment with 3 tasks from the Atari 100k (Bellemare et al., 2013) benchmark. Here, we changed the regular Q-network of the SR-SPR (RR=2) model (D’Oro et al., 2022) to a BroNet, and considered changing the reset schedules to better fit the plasticity of the BroNet model. As depicted in the Table below, applying BroNet to discrete, image-based tasks is a promising avenue for future research.
Table 8: We replace the Q-network in SR-SPR with BroNet, while keeping the standard convolutional encoder. We test two values of reset interval (RI) and shrink-and-perturb (SP) and find that these hyperparameters impact the performance of the BroNet agent. 5 seeds.
Task SR-SPR SR-SPR-BroNet SR-SPR-BroNet SR-SPR-BroNet
(RI=20k;SP=0.8) (RI=20k;SP=0.8) (RI=20k;SP=0.5) (RI=5k;SP=0.8)
Pong -10.5 4.8 10.2 -12.0
Seaquest 714.7 399.0 420.7 782.0
Breakout 24.5 13.5 13.7 33.3
Multi-Task RL
Finally, we evaluate whether the increased model size improves the agent’s capability in a multi-task setup (Yu et al., 2020; Hansen et al., 2023). To this end, we compare the BRO (Fast) to SAC on the MT50 benchmark from MetaWorld, which comprises 50 different tasks. To accommodate the task diversity, we increase the width of all models twofold and pass a one-hot vector representing the task identifier in the network input. Otherwise, both algorithms are run with the same hyperparameters as the main experiments. We present the results in Figure 19.
Figure 19: We compare BRO (Fast) with SAC on the multi-task benchmark. 3 seeds
Appendix F Training Curves
We present the aggregated performance of BRO compared to other baselines at Dog & Humanoid tasks, DeepMind Control Suite, Metaworld and MyoSuite in Figure 20 together with summarized performance results at $100k$, $200k$, $500k$ and $1M$ steps in Table 9. The performance on each of the $40$ individual tasks in shown in Figures 21, 23, 22.
@@ Appendix F Training Curves
Table 9: Summary of IQM results of BRO and other agents evaluated on $40$ tasks from DeepMind Control Suite, Metaworld and MyoSuite achieved at $100k$, $200k$, $500k$ and $1M$ steps. BRO achieves better results than other state-of-the-art agents (both model-based and model-free) while featuring great sample efficiency.